6 (3x-9)= 6+7(8x +4)
Answer by fractalier(6550)

From 6 (3x-9)= 6+7(8x +4), let us distribute and collect like terms...
18x - 54 = 6 + 56x + 28
18x - 54 = 34 + 56x
-88 = 38x
x = -88/38 = -44/19